OpenStudy (anonymous):

The ratio of two numbers is 5 to 4 and their sum is 180. What is their difference ?

OpenStudy (anonymous):

\[=180\div(5+4)\] \[=20\]

OpenStudy (anonymous):

\[=(20\times5)-(20\times4)\] \[=100-80\] \[=?\] solve it @sydneyarnea

OpenStudy (perl):

We are given x/y = 5/4 x + y = 180 solve for x,y then find the difference x - y